# LLM-Driven Composite Neural Architecture Search for Multi-Source RL State Encoding

### v1

@@ Abstract
Designing state encoders for reinforcement learning (RL) with multiple information sources—such as sensor measurements, time-series signals, image observations, and textual instructions—remains underexplored and often requires manual design. We formalize this challenge as a problem of composite neural architecture search (NAS), where multiple source-specific modules and a fusion module are jointly optimized. Existing NAS methods overlook useful side information from the intermediate outputs of these modules—such as their representation quality—limiting sample efficiency in multi-source RL settings. To address this, we propose an LLM-driven NAS pipeline that leverages language-model priors and intermediate-output signals to guide sample-efficient search for high-performing composite state encoders. On a mixed-autonomy traffic control task, our approach discovers higher-performing architectures with fewer candidate evaluations than traditional NAS baselines and the LLM-based GENIUS framework.
@@ 3.2 LLM-Driven Neural Architecture Generation
We begin with an expert-designed initial architecture. At each subsequent iteration, we query an LLM with a textual prompt that summarizes the current set of architecture modules and their associated performance. The LLM then responds with one or a batch of new composite architecture candidates.
For each module (including the source-specific encoders and the fusion block), the search is restricted to a module-specific architecture space typically used for that type of neural network (e.g., CNNs for image inputs, GRUs for textual instructions, transformers for time-series inputs, and FFNs for vector inputs). To illustrate this, Figure˜6 and Table˜1 in Appendix˜B provide an example of a composite architecture for a mixed-autonomy traffic-control benchmark, involving transformers for time-series encoders and FFNs for vector inputs. Similarly, Figure˜9 and Table˜2 provide an example of a composite architecture for the MiniGrid goal-oriented task benchmark, involving an image encoder, a text encoder, and a fusion encoder, along with their corresponding search spaces. Likewise, Figure˜11 provides an example of a composite architecture for the ManiSkill robotic control benchmark, involving an RGB encoder, an information encoder, and a fusion encoder.
@@ 4 Experiment: RL-Based Mixed-Autonomy Traffic Control
We evaluate our method on an RL-based mixed-autonomy traffic control task studied in cheng2025learning in which both connected autonomous vehicles (CAVs) and human-driven vehicles coexist in the same environment. The CAV penetration ratio is set to 0.9. At each environment step, the observation contains three distinct input sources: (i) the temporal traffic evolution of key metrics (e.g., speed, density, and flow rate), (ii) the current traffic state (lane-specific densities, speed distributions, and CAV penetration ratio), and (iii) the distribution history of the vehicle sequence. A schematic of this traffic control scenario is illustrated in Figure˜1(a). The presence of multiple sources of inputs makes this benchmark suitable for evaluating composite state encoders.
@@
Figure˜1(b) shows the average traffic speed as a function of the number of evaluated architecture candidates. Both of our LACER variants significantly outperform the expert-designed architecture, traditional NAS baselines, and the LLM-based GENIUS baseline. These results demonstrate that combining LLM-based priors with composite state encoding and richer performance signals leads to more sample-efficient architecture search in RL settings.
@@ Appendix A Methodology Illustration
Unlike traditional NAS methods for supervised learning (Figure˜2), our approach is designed to optimize multi-source state encoders in RL.
@@
This leads to composite NAS with composite architecture design of the multi-source RL state encoder and corresponding search space (Figure˜3).
@@
Compared to other LLM-based NAS methods (e.g., GENIUS in Figure˜4), our approach is designed to enhance both sample efficiency and solution quality when searching for composite neural architectures for state encoders in RL.
@@
To achieve this, we incorporate evaluation metrics of the initial architecture into the initial prompt to provide richer prior context. Additionally, beyond standard task metric, we introduce two supplementary performance signals—average reward and feature information (i.e., representation quality)—as comprehensive feedback to the LLM, enabling iterative refinement of candidate architectures (Figure˜5).
@@ B.1 RL-based mixed-autonomy traffic control
Following cheng2025learning, the architecture of the state encoder includes a time encoder, traffic encoder, and sequence encoder for respective data encoding while the fusion encoder processes their concatenated outputs to generate the encoded state for RL training, as illustrated in Figure˜6. All modules use Transformer (chosen over recurrent alternatives such as LSTMs, given its superior performance on sequential data) with FFNs. Time and sequence encoders add multi-head self-attention (MHSA) to capture their higher complexity, temporal variability, and dynamics, unlike preprocessed, macro-level, weakly temporal traffic data, and the fusion encoder which focused on integration without extra temporal processing uses only FFN.
@@
Based on the state encoder architecture with multiple modules of distinct functions and specific architecture types in Figure˜6, we define module-specific search spaces, as detailed in Table˜1. Following chitty2022neural, key design choices commonly used in neural architecture search for Transformers are included: hidden layer dimension (denoted as "dimension" in the table), dimension expansion ratio ("ratio"), and number of neural network layers ("depth") for the FFN of each module. For modules with MHSA, the number of attention heads ("heads") is additionally included; for modules with only FFN, the type of activation function ("activation") is included instead. Common value ranges are set for all design choices, resulting in a total of approximately 26 million possible architectures throughout the search space.
@@
RL training requires sufficient steps for policy convergence, typically manifested by reward. In the mixed-autonomy traffic control settings simulated via SUMO, traffic flow arrives with a fixed periodic distribution, causing observed average vehicle speed to exhibit corresponding periodicity. Thus, RL policy evaluation also requires adequate steps to encompass multiple such cycles. To balance RL training convergence, evaluation comprehensiveness, and training cost, we analyzed the average reward and average speed (recorded every 1k steps over 1M steps of baseline RL training) as shown in Figure˜7. The results indicate that the average reward converges around 200k steps, while the average speed exhibits a periodicity of approximately 25k steps. Based on this, each candidate state encoder architecture was evaluated by integrating it into the RL framework for 200k steps of training and 50k steps of evaluation.
@@
For the proposed LACER method (applied to NAS for RL state encoders), LLM type and temperature (a hyperparameter regulating LLM output randomness: higher values enhance diversity, lower values improve determinism) significantly affect task performance. To analyze the impact of design choices, such as the base LLM and its temperature, experiments were conducted on LACER using two representative LLMs (Claude Sonnet 4.0, GPT-4) under temperature configurations of 0.0 and 1.0. As shown in Figure˜8, LACER achieved optimal performance with Claude Sonnet 4.0 (temperature = 1.0), which was thus adopted in the main experiments.
@@ B.2 MiniGrid goal-oriented tasks
Following chevalier2023minigrid, the architecture of the state encoder includes an image encoder and a text encoder for respective data encoding while the fusion encoder processes their concatenated outputs to generate the encoded state for RL training, as illustrated in Figure˜9. The image encoder uses convolution neural networks (CNNs) added with a pooling layer to process image observation. The text encoder uses an embedding layer added with a gate recurrent unit (GRU) to represent the text instruction. The fusion encoder concatenates the processed data of image encoder and text encoder to merge them into final representation as state.
@@
Based on the state encoder architecture with multiple modules of distinct functions and specific architecture types in Figure˜9, we define module-specific search spaces, as detailed in Table˜2. Following chitty2022neural, key hyperparameters commonly used in NAS for CNNs with pooling layers are included: the type of pooling method (denoted as "pooling type" in the table), the layer where the CNN is followed with the pooling layer ("pooling layer"), the type of activation function ("activation"), the size of kernels ("kernel size"), the number of channels ("channel number") and number of neural network layers ("depth") for the image encoder. Following chitty2022neural, key parameters commonly used in NAS for RNNs with embedding layer are included: the dimension of embedding layer ("dimension"), the hidden size of RNNs ("hidden size"), the ratio of drop-out in RNNs ("dropout") and number of neural network layers ("depth") for the text encoder. Following chitty2022neural, key variance of fusion method could be summarized as: the merge type ("merge type"), the dimension of fusion network ("dimension"), the type of activation function ("activation") and the hidden size of networks ("hidden size"). Common value ranges are set for all search parameters, resulting in a total of approximately 19 million possible architectures throughout the search space.
@@
RL training requires sufficient steps for policy convergence, typically manifested by reward. In the MiniGrid goal-oriented tasks, agent uses reshaped return to train policy and requires several trials to finish the task under the trained policy. Thus, RL policy evaluation also requires adequate steps to encompass multiple such trials. To balance RL training convergence, evaluation comprehensiveness, and the cost of evaluating state encoder architectures, we analyzed the average reshaped return of training (recorded every 2,048 steps over 2M steps of baseline RL training) and average return of evaluation (recorded every step over 500 steps of baseline evaluation) as shown in Figure˜10. The results indicate that the average reshaped return converges around 1M steps, while the average return converges around 100 steps. Based on this, each candidate state encoder architecture was evaluated by integrating it into the RL framework for 1M steps of training and 100 steps of evaluation.
@@ B.3 ManiSkill robotic control
Following tao2024maniskill3, the architecture of the state encoder includes an RGB encoder and a information encoder for respective data encoding while the fusion encoder processes their concatenated outputs to generate the encoded state for RL training, as illustrated in Figure˜11. The RGB encoder uses CNNs added with FFNs to process RGB observations. The information encoder uses FFNs to represent other state information. The fusion encoder concatenates the processed data of image encoder and text encoder to merge them into final representation as state.
@@ Appendix C Additional Experiment Results
To verify the necessity of each core module in the proposed LACER method, we designed a series of ablation experiments, in which three key prompt components were removed, respectively: the feature information (FI), the average reward (RI), and the initial architecture evaluation (IE). This resulted in three variant methods: LACER-1 without FI, LACER-1 without FI + RI and LACER-1 without FI + RI + IE. The results in Fig.˜12 indicate that the original LACER-1 method achieves the best performance. When any of the components mentioned above is removed, the performance of LACER deteriorates significantly. This phenomenon demonstrates that each of these components in our LACER method is useful and essential, and their collaborative operation contributes to the superior performance of the proposed method.
@@
The time cost of traditional NAS methods is composed of the evaluation time of each candidate and the search time of each generation, while the search cost of LLM-based NAS methods includes the query time of LLM and the processing time of analyzing results of candidates and constructing prompts. To analyze the impact of query time, we run LACER and other baselines on an Intel Core i7 CPU and an NIDIA RTX 4070 GPU to avoid the time cost of node scheduling. We then compare the time cost of the above compositions as shown in Figure˜13, where evaluation refers to the evaluation time, search refers to the search time that does not include query time, and query refers to the query time. The result indicates that the query time of LLM has a negligible impact on the overall time cost compared to the evaluation time.

### v2

@@ Abstract
Designing state encoders for reinforcement learning (RL) with multiple information sources—such as sensor measurements, time-series signals, image observations, and textual instructions—remains underexplored and often requires manual design. We formalize this challenge as a problem of composite neural architecture search (NAS), where multiple source-specific modules and a fusion module are jointly optimized. Existing NAS methods overlook useful side information from the intermediate outputs of these modules—such as their representation quality—limiting sample efficiency in multi-source RL settings. To address this, we propose an LLM-driven NAS pipeline in which the LLM serves as a neural architecture design agent, leveraging language-model priors and intermediate-output signals to guide sample-efficient search for high-performing composite state encoders. On a mixed-autonomy traffic control task, our approach discovers higher-performing architectures with fewer candidate evaluations than traditional NAS baselines and the LLM-based GENIUS framework.
@@ 3 LACER: An LLM → State Encoder → RL Pipeline
Figure 1: Our LLM design agent → architecture generation → RL training & evaluation pipeline. In this pipeline, the LLM serves as a neural architecture design agent that perceives performance signals, reasons over conversation history, and decides on candidate architectures, facilitating iterative refinement via closed-loop interaction.
@@ 3.2 LLM-Driven Neural Architecture Generation
We begin with an expert-designed initial architecture. At each subsequent iteration, we query the LLM with a concise summary of the current architecture modules and their associated performance. The LLM then produces one or a batch of new composite architecture candidates. In this pipeline, the LLM functions as a neural architecture design agent, generating new module configurations informed by the feedback signals from previously evaluated candidates.
For each module (including the source-specific encoders and the fusion block), the search is restricted to a module-specific architecture space typically used for that type of neural network (e.g., CNNs for image inputs, GRUs for textual instructions, transformers for time-series inputs, and FFNs for vector inputs). To illustrate this, Figure˜7 and Table˜1 in Appendix˜B provide an example of a composite architecture for a mixed-autonomy traffic-control benchmark, involving transformers for time-series encoders and FFNs for vector inputs. Similarly, Figure˜10 and Table˜2 provide an example of a composite architecture for the MiniGrid goal-oriented task benchmark, involving an image encoder, a text encoder, and a fusion encoder, along with their corresponding search spaces. Likewise, Figure˜12 provides an example of a composite architecture for the ManiSkill robotic control benchmark, involving an RGB encoder, an information encoder, and a fusion encoder.
@@ 4 Experiment: RL-Based Mixed-Autonomy Traffic Control
We evaluate our method on an RL-based mixed-autonomy traffic control task studied in cheng2025learning in which both connected autonomous vehicles (CAVs) and human-driven vehicles coexist in the same environment. The CAV penetration ratio is set to 0.9. At each environment step, the observation contains three distinct input sources: (i) the temporal traffic evolution of key metrics (e.g., speed, density, and flow rate), (ii) the current traffic state (lane-specific densities, speed distributions, and CAV penetration ratio), and (iii) the distribution history of the vehicle sequence. A schematic of this traffic control scenario is illustrated in Figure˜2(a). The presence of multiple sources of inputs makes this benchmark suitable for evaluating composite state encoders.
@@
Figure˜2(b) shows the average traffic speed as a function of the number of evaluated architecture candidates. Both of our LACER variants significantly outperform the expert-designed architecture, traditional NAS baselines, and the LLM-based GENIUS baseline. These results demonstrate that combining LLM-based priors with composite state encoding and richer performance signals leads to more sample-efficient architecture search in RL settings.
@@ Appendix A Methodology Illustration
Unlike traditional NAS methods for supervised learning (Figure˜3), our approach is designed to optimize multi-source state encoders in RL.
@@
This leads to composite NAS with composite architecture design of the multi-source RL state encoder and corresponding search space (Figure˜4).
@@
Compared to other LLM-based NAS methods (e.g., GENIUS in Figure˜5), our approach is designed to enhance both sample efficiency and solution quality when searching for composite neural architectures for state encoders in RL.
@@
To achieve this, we incorporate evaluation metrics of the initial architecture into the initial prompt to provide richer prior context. Additionally, beyond standard task metric, we introduce two supplementary performance signals—average reward and feature information (i.e., representation quality)—as comprehensive feedback to the LLM, enabling iterative refinement of candidate architectures (Figure˜6).
@@ B.1 RL-based mixed-autonomy traffic control
Following cheng2025learning, the architecture of the state encoder includes a time encoder, traffic encoder, and sequence encoder for respective data encoding while the fusion encoder processes their concatenated outputs to generate the encoded state for RL training, as illustrated in Figure˜7. All modules use Transformer (chosen over recurrent alternatives such as LSTMs, given its superior performance on sequential data) with FFNs. Time and sequence encoders add multi-head self-attention (MHSA) to capture their higher complexity, temporal variability, and dynamics, unlike preprocessed, macro-level, weakly temporal traffic data, and the fusion encoder which focused on integration without extra temporal processing uses only FFN.
@@
Based on the state encoder architecture with multiple modules of distinct functions and specific architecture types in Figure˜7, we define module-specific search spaces, as detailed in Table˜1. Following chitty2022neural, key design choices commonly used in neural architecture search for Transformers are included: hidden layer dimension (denoted as "dimension" in the table), dimension expansion ratio ("ratio"), and number of neural network layers ("depth") for the FFN of each module. For modules with MHSA, the number of attention heads ("heads") is additionally included; for modules with only FFN, the type of activation function ("activation") is included instead. Common value ranges are set for all design choices, resulting in a total of approximately 26 million possible architectures throughout the search space.
@@
RL training requires sufficient steps for policy convergence, typically manifested by reward. In the mixed-autonomy traffic control settings simulated via SUMO, traffic flow arrives with a fixed periodic distribution, causing observed average vehicle speed to exhibit corresponding periodicity. Thus, RL policy evaluation also requires adequate steps to encompass multiple such cycles. To balance RL training convergence, evaluation comprehensiveness, and training cost, we analyzed the average reward and average speed (recorded every 1k steps over 1M steps of baseline RL training) as shown in Figure˜8. The results indicate that the average reward converges around 200k steps, while the average speed exhibits a periodicity of approximately 25k steps. Based on this, each candidate state encoder architecture was evaluated by integrating it into the RL framework for 200k steps of training and 50k steps of evaluation.
@@
For the proposed LACER method (applied to NAS for RL state encoders), LLM type and temperature (a hyperparameter regulating LLM output randomness: higher values enhance diversity, lower values improve determinism) significantly affect task performance. To analyze the impact of design choices, such as the base LLM and its temperature, experiments were conducted on LACER using two representative LLMs (Claude Sonnet 4.0, GPT-4) under temperature configurations of 0.0 and 1.0. As shown in Figure˜9, LACER achieved optimal performance with Claude Sonnet 4.0 (temperature = 1.0), which was thus adopted in the main experiments.
@@ B.2 MiniGrid goal-oriented tasks
Following chevalier2023minigrid, the architecture of the state encoder includes an image encoder and a text encoder for respective data encoding while the fusion encoder processes their concatenated outputs to generate the encoded state for RL training, as illustrated in Figure˜10. The image encoder uses convolution neural networks (CNNs) added with a pooling layer to process image observation. The text encoder uses an embedding layer added with a gate recurrent unit (GRU) to represent the text instruction. The fusion encoder concatenates the processed data of image encoder and text encoder to merge them into final representation as state.
@@
Based on the state encoder architecture with multiple modules of distinct functions and specific architecture types in Figure˜10, we define module-specific search spaces, as detailed in Table˜2. Following chitty2022neural, key hyperparameters commonly used in NAS for CNNs with pooling layers are included: the type of pooling method (denoted as "pooling type" in the table), the layer where the CNN is followed with the pooling layer ("pooling layer"), the type of activation function ("activation"), the size of kernels ("kernel size"), the number of channels ("channel number") and number of neural network layers ("depth") for the image encoder. Following chitty2022neural, key parameters commonly used in NAS for RNNs with embedding layer are included: the dimension of embedding layer ("dimension"), the hidden size of RNNs ("hidden size"), the ratio of drop-out in RNNs ("dropout") and number of neural network layers ("depth") for the text encoder. Following chitty2022neural, key variance of fusion method could be summarized as: the merge type ("merge type"), the dimension of fusion network ("dimension"), the type of activation function ("activation") and the hidden size of networks ("hidden size"). Common value ranges are set for all search parameters, resulting in a total of approximately 19 million possible architectures throughout the search space.
@@
RL training requires sufficient steps for policy convergence, typically manifested by reward. In the MiniGrid goal-oriented tasks, agent uses reshaped return to train policy and requires several trials to finish the task under the trained policy. Thus, RL policy evaluation also requires adequate steps to encompass multiple such trials. To balance RL training convergence, evaluation comprehensiveness, and the cost of evaluating state encoder architectures, we analyzed the average reshaped return of training (recorded every 2,048 steps over 2M steps of baseline RL training) and average return of evaluation (recorded every step over 500 steps of baseline evaluation) as shown in Figure˜11. The results indicate that the average reshaped return converges around 1M steps, while the average return converges around 100 steps. Based on this, each candidate state encoder architecture was evaluated by integrating it into the RL framework for 1M steps of training and 100 steps of evaluation.
@@ B.3 ManiSkill robotic control
Following tao2024maniskill3, the architecture of the state encoder includes an RGB encoder and a information encoder for respective data encoding while the fusion encoder processes their concatenated outputs to generate the encoded state for RL training, as illustrated in Figure˜12. The RGB encoder uses CNNs added with FFNs to process RGB observations. The information encoder uses FFNs to represent other state information. The fusion encoder concatenates the processed data of image encoder and text encoder to merge them into final representation as state.
@@ Appendix C Additional Experiment Results
To verify the necessity of each core module in the proposed LACER method, we designed a series of ablation experiments, in which three key prompt components were removed, respectively: the feature information (FI), the average reward (RI), and the initial architecture evaluation (IE). This resulted in three variant methods: LACER-1 without FI, LACER-1 without FI + RI and LACER-1 without FI + RI + IE. The results in Fig.˜13 indicate that the original LACER-1 method achieves the best performance. When any of the components mentioned above is removed, the performance of LACER deteriorates significantly. This phenomenon demonstrates that each of these components in our LACER method is useful and essential, and their collaborative operation contributes to the superior performance of the proposed method.
@@
The time cost of traditional NAS methods is composed of the evaluation time of each candidate and the search time of each generation, while the search cost of LLM-based NAS methods includes the query time of LLM and the processing time of analyzing results of candidates and constructing prompts. To analyze the impact of query time, we run LACER and other baselines on an Intel Core i7 CPU and an NIDIA RTX 4070 GPU to avoid the time cost of node scheduling. We then compare the time cost of the above compositions as shown in Figure˜14, where evaluation refers to the evaluation time, search refers to the search time that does not include query time, and query refers to the query time. The result indicates that the query time of LLM has a negligible impact on the overall time cost compared to the evaluation time.